Amazon has issued 22-year semi-annual coupon bonds that have seven years left until maturity. If the coupon rate is 6.70%, and the YTM is 6.80%, what is the current bond price?
Answer #1

Assuming face value to be $1000

Number of periods = 7 * 2 = 14

Coupon = (0.067 * 1000) / 2 = 33.5

Rate = 6.80% / 2 = 3.4%

Current bond price = Coupon * [1 - 1 / (1 + r)n] / r + FV / (1 + r)n

Current bond price = 33.5 * [1 - 1 / (1 + 0.034)14] / 0.034 + 1000 / (1 + 0.034)14

Current bond price = 33.5 * [1 - 0.626199] / 0.034 + 626.199126

Current bond price = 33.5 * 10.994147 + 626.199126

Current bond price = $994.50